# Lattice Protein Folding with Variational Annealing

### v1

@@ 4 Conclusion
In this paper, we introduce a novel upper-bound training scheme with masking to find the lowest energy fold of the 2D HP lattice protein folding. Using this scheme and by supplementing Dilated RNNs with annealing through temperature-like fluctuations, we find that our method can find the optimal folds of prototypical lattice folding benchmarks with system sizes up to 60 beads. We demonstrate that it is possible to mask moves that lead to invalid folds without compromising the autoregressive sampling feature of RNNs. We also devise a free energy upper bound loss function that enhances the trainability of RNNs that are prone to get compromised by masking the RNN probabilities.
Our scheme is generalizable to three spatial dimensions and also other lattice protein folding models with more than two alphabets such as the 20-letter Miyazawa-Jernigan model [42] by enlarging the sizes of the one-hot inputs $\bm{x}_{i}$ without compromising inference speed. We also expect inference time to be reduced by introducing an encoder network to enable just-in-time inference of low-energy folds [15]. For a broader scope, we also believe that our scheme could lead to a promising machine learning-based solution to a wide class of constrained combinatorial optimization problems.
@@ Acknowledgments
EMI acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Economic Development, Job Creation and Trade. We also want to thank Caleb Schultz Kisby for holding valuable discussions about the project. Computer simulations were made possible thanks to the Digital Research Alliance of Canada cluster.
@@ Appendix A Hyperparameters
Table 3: In (a), we introduce an additional hyperparameter, $N_{\text{warmup}}$, which is used to perform an optional ‘warm start’ before the annealed training process. In this warm start, we perform exactly $N_{\text{warmup}}$ gradient descent steps at fixed temperature $T_{0}$. Additionally, note that $T_{0}$ is the initial temperature, $M$ is the number of samples used for each training step, $N_{\rm{train}}$ is the number of training steps during each annealing step, and $N_{\rm annealing}$ is the total number of annealing steps. In (b), we highlight that we use the same hidden unit size for each RNN layer.

### v2

@@ 4 Conclusion
In this paper, we introduce a novel upper-bound training scheme with masking to find the lowest energy fold of the 2D HP lattice protein folding. Using this scheme and by supplementing Dilated RNNs with annealing through temperature-like fluctuations, we find that our method can find the optimal folds of prototypical lattice folding benchmarks with system sizes up to 60 beads. We demonstrate that it is possible to mask moves that lead to invalid folds without compromising the autoregressive sampling feature of RNNs. We also devise a free energy upper bound loss function that enhances the trainability of RNNs that are prone to get compromised by masking the RNN probabilities. We note that there is still room for exploring variational annealing with other neural network architectures, such as Transformers [25] on the lattice folding problem and beyond.
To motivate future lines of inquiry, we want to highlight that our scheme is generalizable to folding HP sequences in three spatial dimensions with no significant increase in computational complexity in the forward pass with respect to input size $N$. The 3D model introduces two additional moves to the base set of four moves, which are the ‘forward’ and ‘backward’ moves on the $z$-axis. This setting can be easily implemented by increasing the size of RNN output vector $\boldsymbol{d}_{i}$ to six and the input vector $\bm{x}_{i}$ to eight. The energy function will then take into account six neighbors instead of four. Our method can also generalize to other lattice protein folding models with more than two alphabets such as the 20-letter Miyazawa-Jernigan model [36] by enlarging the sizes of the one-hot inputs $\bm{x}_{i}$ without compromising inference speed. Although our scheme extends implementation to 3D and larger-alphabet models, varying model dynamics may pose unique learning challenges. Investigating our method’s interplay with a family of close-knit folding problems will help distinguish generalizable principles from those needing tailored adjustments.
We also expect inference time to be reduced by introducing an encoder network to enable just-in-time inference of low-energy folds [13]. Lastly, for a broader scope, we also believe that our scheme could lead to a promising machine learning-based solution to a wide class of constrained combinatorial optimization problems such as graph coloring and the traveling salesman problem.
@@ Acknowledgments
EMI and MH acknowledge support from the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Economic Development, Job Creation and Trade. SAK also wants to thank Caleb Schultz Kisby for holding valuable discussions about the project. Computer simulations were made possible thanks to the Digital Research Alliance of Canada cluster.
Code availability
Our code is publicly available on GitHub at https://github.com/shoummoak/VA-HP-Lattice.
@@ Appendix A Hyperparameters
Table 3: In (a), we introduce an additional hyperparameter $N_{\text{warmup}}$, which is used to perform an optional ‘warm start’ before the annealed training process. In this warm start, we perform exactly $N_{\text{warmup}}$ gradient descent steps at fixed temperature $T_{0}$. Additionally, note that $T_{0}$ is the initial temperature, $M$ is the number of samples used for each training step, $N_{\rm{train}}$ is the number of training steps during each annealing step, and $N_{\rm annealing}$ is the total number of annealing steps. In (b), we highlight that we use the same hidden unit size for each RNN layer.